what's next 5 numbers in this sequence? 1, 6, 12, 17, 34, 39, 78...
Answer by KMST(5328)

83, 166, 171, 342, 347
Adding 11 to the 1st and 2nd numbers you get the 3rd and 4th respectively.
1+11=12, 6+11=17
Adding 22 to the 3rd and 4th numbers you get the 5th and 6th respectively.
12+22=34, 17+22=39
Adding 44 to the 5th and 6th numbers you get the 7th and 8th respectively.
34+44=78, 39+44=83
Adding 88 to the 7th and 8th numbers you get the 9th and 10th respectively.
78+88=166, 83+88=171
Adding 176 to the 9th and 10th numbers you get the 11th and 12th respectively.
166+176=342, 171+176=347